OpenStudy (anonymous):

A 12.0 μF capacitor is charged to a potential of 50.0 V and discharged through a 175 Ω resistor. How long does it take the capacitor to lose half its charge? A 12.0 μF capacitor is charged to a potential of 50.0 V and discharged through a 175 Ω resistor. How long does it take the capacitor to lose half its charge? @Physics

OpenStudy (jamesj):

I'd be interested to see the answer to this.

OpenStudy (anonymous):

the answer is 1.5*10^-3 seconds, but I don't know how to calculate it, it involves the equation for a charging capacitor. \[\xi/R*e^(-t/RC)\]

OpenStudy (anonymous):

The function to calculated voltage is: \[\Large {V_c}\left( t \right) = {V_i}{e^{ - \frac{t}{{RC}}}}\] You can work backwards from this to find the time. Step by step: \[\Large \begin{array}{l} {V_c}\left( t \right) = {V_i}{e^{ - \frac{t}{{RC}}}}\\ \ln \left( {{V_c}} \right) = \ln \left( {{V_i}{e^{ - \frac{t}{{RC}}}}} \right) = \ln \left( {{V_i}} \right) + \ln \left( {{e^{ - \frac{t}{{RC}}}}} \right)\\ \ln ({V_c}) = \ln ({V_i}) - \frac{t}{{RC}}\\ \ln ({V_c}) - \ln ({V_i}) = - \frac{t}{{RC}}\\ \ln \left( {\frac{{{V_c}}}{{{V_i}}}} \right) = - \frac{t}{{RC}}\\ t = - RC \cdot \ln \left( {\frac{{{V_c}}}{{{V_i}}}} \right)\\ t = - 175 \cdot \left( {12 \times {{10}^{ - 6}}} \right) \cdot \ln \left( {\frac{{25}}{{50}}} \right)\\ = 1.4556 \times {10^{ - 3}}{\rm{s}} \end{array}\] A handy formula to keep around would be: \[\Large t = - RC \cdot \ln \left( p \right)\]Where p is the decimal percentage of voltage change (i.e. Vc/Vi).